# Exploring Feedback Generation in Automated Skeletal Movement Assessment: A Comprehensive Overview

### v1

@@
Abstract—The application of machine-learning solutions to movement assessment from skeleton videos has attracted significant research attention in recent years. This advancement has made rehabilitation at home more accessible, utilizing movement assessment algorithms that can operate on affordable equipment for human pose detection from 2D or 3D videos. While the primary objective of automatic assessment tasks is to score movements, the automatic generation of feedback highlighting key movement issues has the potential to significantly enhance and accelerate the rehabilitation process. In this study, we explain the types of feedback that can be generated, review existing solutions for automatic feedback generation, and discuss future research directions. To our knowledge, this is the first comprehensive review of feedback generation in skeletal movement assessment.
@@ 1 Introduction
Research into analyzing human movements through skeletal videos has seen significant growth in recent years [22]. Such analysis serves various purposes, including action recognition [37], person identification [27], prediction of medical conditions [33, 16], and movement assessment [19, 41, 17, 7, 34, 8]. Importantly, these analyses often play a crucial role in monitoring and supporting rehabilitation processes [20, 26, 14, 15].
@@
The primary drawback of automatic assessment at home lies in providing feedback and guidance for improvement. Similar to how loss-gradient guides iterative optimization algorithms by indicating direction, feedback can enhance performance over time. In face-to-face assessments, professionals can offer direct guidance to patients between or even during performances. However, in an automatic assessment system, providing such guidance is not as straightforward or immediate. It may require the utilization of advanced machine-learning techniques and human-computer interaction practices.
@@ 2 Movement Assessment
Movement assessment, as one of the applications of movement analysis from skeletal videos, presents a relatively challenging task. Unlike action recognition, which typically involves classifying movements from a high-level temporal and skeletal perspective, movement assessment requires recognizing low-level details crucial for performance quality and closely attending to them at specific stages of movement progress [19]. Automatic movement assessment solutions find application across various domains, including rehabilitative medical assessments [43, 20, 15, 33, 26, 38, 4, 45, 5, 47, 31, 11, 28], gymnastics [40, 3, 24, 32], physical skills [13, 23, 46], and Olympic sports [29].
@@
Feedback generation, which holds the potential to accelerate movement learning and, consequently, the rehabilitation process [19, 17, 8], is occasionally addressed by studies focusing on movement assessment [20, 26, 28, 40, 24, 30]. Feedback is sometimes equated with interpretability, which entails understanding why a specific outcome is observed how input changes will alter it [17]. As we discuss in this paper, the concept of feedback generation can be interpreted in multiple ways. Evaluating the quality of generated feedback can be challenging. From a data perspective, annotating training and test samples with their expected feedback may be considerably more complex than solely annotating them with their expected performance scores. Moreover, in certain scenarios, multiple feedback outputs may be deemed correct or partially correct, posing challenges for evaluation. In the subsequent sections of this paper, we delve deeper into feedback generation.
@@ 3 Feedback Generation
Defining ”feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally, textually, or visually [1, 2], and occasionally through physical means [18].
@@
Feedback Generation Mode.
In the domain of automatic movement assessment, movements are represented as sequences of skeleton poses, inherently possessing a temporal dimension. Recognizing the stages of movement progression during assessment, termed as temporal alignment, is crucial due to variations in movement progress across different performances [19]. Some studies employ online alignment techniques using tools like Hidden Markov Models (HMMs) [39, 9] and Recurrent Neural Networks (RNNs) [25, 4, 35, 36, 40], while others opt for offline methods such as Dynamic Time Warping (DTW) [6, 43, 12, 13, 46], piecewise warping between temporal points-of-interest [20], and DTAN [42, 26]. However, offline temporal alignment poses a challenge for generating online feedback during performance analysis, as it only initiates after the entire performance has concluded. Conversely, online temporal alignment allows for ongoing analysis after each frame input, potentially enabling the production of online feedback. This capability holds promise in bridging one of the significant gaps between automatic and frontal assessments.
@@
Creating textual feedback that is both intuitive and informative presents challenges, as discussed earlier. In contrast, visual feedback holds the potential to be more intuitive and informative, and sometimes even easier to generate. For instance, presenting multiple first-order unbound feedback labels simultaneously as an animation on a skeleton figure inherently incorporates skeletal and temporal aggregation, as illustrated in Figure 3.
Figure 3: A proposed method for offering highly relevant and intuitive feedback by visualizing only the first-order deviations together. In this example, although the three reported deviations are first-order, which could be overwhelming in text form, the visualization clearly indicates that the main issue is the right elbow being too low. In other words, skeletal aggregation occurs naturally through perception.

### v2

@@
Abstract—The application of machine-learning solutions to movement assessment from skeleton videos has attracted significant research attention in recent years. This advancement has made rehabilitation at home more accessible, utilizing movement assessment algorithms that can operate on affordable equipment for human pose detection and analysis from 2D or 3D videos. While the primary objective of automatic assessment tasks is to score movements, the automatic generation of feedback highlighting key movement issues has the potential to significantly enhance and accelerate the rehabilitation process. While numerous research works exist in the field of automatic movement assessment, only a handful address feedback generation. In this study, we explain the types of feedback that can be generated, review existing solutions for automatic feedback generation, and discuss future research directions. To our knowledge, this is the first comprehensive review of feedback generation in skeletal movement assessment.
@@ 1 Introduction
Research into analyzing human movements through skeletal videos has seen significant growth in recent years [24]. Such analysis serves various purposes, including action recognition [40], person identification [30], prediction of medical conditions [36, 18], and movement assessment [44, 19, 8, 37, 10, 21]. Importantly, these analyses often play a crucial role in monitoring and supporting rehabilitation processes [9, 22, 29, 16, 17].
@@
The primary drawback of automatic assessment at home lies in providing feedback and guidance for improvement. Similar to how loss-gradient guides iterative optimization algorithms by indicating direction, feedback plays a crucial role in guiding patients towards improved performance over time. In face-to-face assessments, professionals can offer direct guidance to patients between or even during performances. However, in an automatic assessment system, providing such guidance is not as straightforward or immediate. It may require the utilization of advanced machine-learning techniques and human-computer interaction practices.
@@ 2 Movement Assessment
Movement assessment, as one of the applications of movement analysis from skeletal videos, presents a relatively challenging task. Unlike action recognition, which typically involves classifying movements from a high-level temporal and skeletal perspective, movement assessment requires recognizing low-level details crucial for performance quality and closely attending to them at specific stages of movement progress [21]. Automatic movement assessment solutions find application across various domains, including rehabilitative medical assessments [46, 22, 17, 36, 29, 41, 5, 48, 6, 50, 34, 13, 31, 28], gymnastics [43, 4, 26, 35], physical skills [15, 25, 49], and Olympic sports [32].
@@
Feedback generation, which holds the potential to accelerate movement learning and, consequently, the rehabilitation process [44, 19, 21, 10], is occasionally addressed by studies focusing on movement assessment [3, 22, 29, 31, 43, 26, 33]. Feedback is sometimes equated with interpretability, which entails understanding why a specific outcome is observed how input changes will alter it [19]. As we discuss in this paper, the concept of feedback generation can be interpreted in multiple ways. Evaluating the quality of generated feedback can be challenging. From a data perspective, annotating training and test samples with their expected feedback may be considerably more complex than solely annotating them with their expected performance scores. Moreover, in certain scenarios, multiple feedback outputs may be deemed correct or partially correct, posing challenges for evaluation. In the subsequent sections of this paper, we delve deeper into feedback generation.
@@ 3 Feedback Generation
Defining ”feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally [3, 26, 31], textually [1, 22, 31, 29], or visually [3, 1, 11, 43, 26], and occasionally through physical means [20]. Table 1 provides a summary of feedback generation solutions, based on the criteria proposed in the following sections.
@@
Offline/Online Feedback Generation Mode.
In the domain of automatic movement assessment, movements are represented as sequences of skeleton poses, inherently possessing a temporal dimension. Recognizing the stages of movement progression during assessment, termed as temporal alignment, is crucial due to variations in movement progress across different performances [21]. Some studies employ online alignment techniques using tools like Hidden Markov Models (HMMs) [42, 11] and Recurrent Neural Networks (RNNs) [27, 5, 38, 39, 43], while others opt for offline methods such as Dynamic Time Warping (DTW) [7, 46, 14, 15, 49], piecewise warping between temporal points-of-interest [22], and DTAN [45, 29]. Offline temporal alignment poses a challenge for generating online feedback during performance analysis, as it only initiates after the entire performance has concluded. Conversely, online temporal alignment allows for ongoing analysis after each frame input, potentially enabling the production of online feedback. This capability holds promise in bridging one of the significant gaps between automatic and frontal assessments.
@@
Creating textual or verbal feedback that is both intuitive and informative presents challenges, as discussed earlier. In contrast, visual feedback holds the potential to be more intuitive and informative, and sometimes even easier to generate. For instance, presenting multiple first-order unbound feedback labels simultaneously as an animation on a skeleton figure inherently incorporates skeletal and temporal aggregation, as illustrated in Figure 3. As previously discussed, several works have proposed solutions for visual feedback generation.
Figure 3: A proposed method for offering highly relevant and intuitive feedback by visualizing only the first-order deviations together. In this example, although the three reported deviations are first-order, which could be overwhelming in textual form, the visualization clearly indicates that the main issue is the right elbow being too low. In other words, skeletal aggregation occurs naturally through perception.
@@ 4 Evaluating Generated Feedback
Ref. Year Movement Domain Modalities Boundness, Order Gen. Mode Temp. Granul. Evaluation Type
[3] 2015 Physical Rehabilitation Visual,Verbal Bound Online Local N/A
[11] 2016 Gait Visual Unbound Online Local Qualitative
[43] 2016 Powerlifting Visual Unbound Online Local Quantitative(Q)
[1, 2] 2016-17 Stroke Rehabilitation Text,Visual Unbound (order=3) Offline Local,Global Qualitative
[26] 2018 Squats and Tai Chi Visual,Verbal Unbound Online Local N/A
[22, 23] 2019 Stroke Rehabilitation Text Unbound (order=2) Offline Local,Global Qualitative
[31] 2020 Stroke Rehabilitation Text,Verbal Unbound (order=N/A) Online N/A N/A
[29] 2023 Hip-Replacement Recovery Text Bound Offline Global Quantitative
Table 1: Existing works dealing with feedback generation. In Evaluation Type, (Q) stands for quantized (bound) feedback labels. Feedback order is only reported for unbound feedback in textual modality.
@@ 5 Discussion
Another aspect we have explored is the feedback modality. We have proposed that, especially in cases where feedback is unbound, generating textual feedback without overwhelming the recipient might be challenging. In contrast, when presented visually, even raw, first-order feedback may be interpreted more intuitively.